OpenStudy (anonymous):

Tristan walks 3/4 of a mile to school every day. He wants to find how long it will take him to walk 6 miles. Use the model to find 6 ÷3/4 , the number of days it will take Tristan to walk 6 miles.

OpenStudy (anonymous):

dividing by three-fourths is the same as multiplying by four-thirds

OpenStudy (anonymous):

so its four- thirds

OpenStudy (anonymous):

no it is six times four thirds

OpenStudy (anonymous):

\[6\div \frac{3}{4}=6\times \frac{4}{3}=2\times 4=8\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

or if you prefer, \[6\div \frac{3}{4}=6\times \frac{4}{3}=\frac{24}{3}=8\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks